OpenStudy (anonymous):

-2/5 +t=5/12 the choices are 49/60 or 7/17

OpenStudy (anonymous):

Can you get the numbers all on one side, then get a common denominator?

OpenStudy (anonymous):

7/17 ??

OpenStudy (anonymous):

Hellooo

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

nooo it's not..

OpenStudy (anonymous):

o wait, why did you...oh nvm. t = 5/12 + 2/5 t = 25/60 + 12/60 t = 37/60

OpenStudy (anonymous):

24/60* thus 49/60